OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24

OpenStudy (lgbasallote):

hint: use the equation \[\large y = kx\]

OpenStudy (anonymous):

y = kx y=12,x=16 so 12 = k*16 k = 3/4 so y = 3x/4 when y=24 24 = 3x/4 so x = 96/3 = 32 x=32

OpenStudy (anonymous):

thank u so much